(x+y-1)dx+(2x+2y+1)dy=0


1
Expert's answer
2021-03-15T17:05:17-0400

We have,


-"(x + y -1)dx = (2x + 2y +1)dy"


"\dfrac{dy}{dx} = -\dfrac{(x+y-1)}{(2x+2y+1)}"


Let "x+y = t"


Differentiating with respect to 'x'


"1 + \dfrac{dy}{dx} = \dfrac{dt}{dx}"


"\dfrac{-(t-1)}{2t+1} = \dfrac{dt}{dx} - 1"


"1 - \dfrac{(t -1)}{(2t+1)} = \dfrac{dt}{dx}"


"\dfrac{dt}{dx} = \dfrac{(2t+1-t+1)}{(2t+1)}"


"\dfrac{(t+2)}{(2t+1)} = \dfrac{t+2}{2(t+\dfrac{1}{2})}"


"2\dfrac{dt}{dx} = \dfrac{t+2}{t+\dfrac{1}{2}}"


"\int\dfrac{(2t+1)}{t+2}dt = \int dx"


"\int \dfrac{(2t+4-3)}{(t+2)}dt = \int dx"


"\int 2dt -3 \int \dfrac{dt}{t+2} = \int dx"


"2t - 3ln|t+2| = x+c"


"2(x+y) - 3ln|x+y+2| = x+c"